A three-cell flashlight draws a current of 0.60A. What is the operating resistance of the light bulb if each cell provides a potential of 1.0V when delivering this current?
Answer #1

here,

potential of each cell , V = 1 v

as there are three cell in series


Vnet = 3*V

vnet = 3 V

current , I = 0.6 A

the operating resistance of the light bulb , R = V/I

R = 3/0.6

R = 5 ohm

the operating resistance of the light bulb is 5 ohm